OpenStudy (moses238):

(6x + 6)/(x^2 - 1) * (x - 1)/(8x^2 - 24x)÷(24)/(x^2 - 9) Simplify

OpenStudy (will.h):

Can you post a screen shot. It's hard to read all that symbols

OpenStudy (moses238):

yes

OpenStudy (moses238):

i tried so many time i just can't get it

OpenStudy (will.h):

Very well. What do you have so far. Did you use factoring or distributive property and divisions?

OpenStudy (moses238):

i dont know what to do first

OpenStudy (will.h):

What is your course and what's the name of the lesson so I can find appropriate way to help you because there are many ways to simplify that.

OpenStudy (moses238):

this is my multiple choice

OpenStudy (angelinaxoxo1):

https://mathway.com/ Try this Site

OpenStudy (moses238):

Rational equations is my lesson

OpenStudy (moses238):

i used that website and it gave me the answer but no explanation

OpenStudy (will.h):

That's more confusing than I thought. But did you try to solve the options and see if the result congruent to the expression than the answer would be right. So far am still working on it but it's really confusing.

OpenStudy (campbell_st):

the basic thing you need to do is factor the quadratics and then factor the linear expressions. remember, the last part is dividing by a fraction so flip and multiply so the expression becomes \[\frac{6(x +1)}{(x+1)(x-1)} \times \frac{(x -1)}{8x(x - 3)} \times \frac{(x -3)(x +3)}{24}\] so remove the common factors and see what you have left